# CLQCD Collaboration

### v2

@@ Abstract
We report a high precision calculation of the isospin vector charge $g_{S,T}$ of the nucleon using recently proposed “blending” method which provides a high-precision stochastic estimate of the all-to-all fermion propagator. By combining the current-involved interpolation operator, which can efficiently cancel the major excited state contaminations, we can extract high-precision $g_{S}$ and $g_{T}$ directly at the physical pion mass. Using 15 $N_{f}=2+1$ lattice ensembles which cover 5 lattice spacing, 5 combinations with the same quark masses and lattice spacing but multiple volumes, including three at the physical pion mass, we report so far most precise lattice QCD prediction $g_{T}^{\rm QCD}=1.0256[78]_{\rm tot}(58)_{\rm stat}(17)_{a}(44)_{\rm FV}(01)_{\chi}(22)_{\rm ex}(05)_{\rm re}$ and $g_{S}^{\rm QCD}=1.107[46]_{\rm tot}(32)_{\rm stat}(04)_{a}(29)_{\rm FV}(01)_{\chi}(13)_{\rm ex}(08)_{\rm re}$ at $\overline{\mathrm{MS}}$ 2 GeV, with the systematic uncertainties from continuum, infinite volume, chiral extrapolations, excited state contamination and also renormalization.
@@
The nucleon iso-vector charges $\bar{u}\Gamma_{X}u-\bar{d}\Gamma_{X}d$—specifically for the axial-vector ($\Gamma_{A}=\gamma_{5}\gamma_{i}$), scalar ($\Gamma_{S}=\mathbf{1}$), and tensor ($\Gamma_{T}=\sigma_{ij}$) currents—are among the simplest matrix elements to compute. This simplicity arises because the disconnected quark diagrams cancel out, up to isospin-breaking (ISB) effects. While the axial current can be tightly constrained from the neutron weak decay, modulo ISB and QED corrections ParticleDataGroup:2024cfk, the scalar and tensor charges are in higher demand. They are particularly crucial for probing TeV-scale new physics Bhattacharya:2011qm; Bhattacharya:2016zcn, enabling high-precision calculations of the proton-neutron mass difference BMW:2014pzb, and providing constraints on parton distribution functions Lin:2017stx; LatticeParton:2022xsd.
However, precise predictions of these charges remain highly challenging. This is primarily due to the competing difficulties of significant excited-state contamination (ESC) at small source-sink separations ($t_{f}$) and exponentially growing statistical noise at large $t_{f}$. Consequently, although sophisticated methods (e.g., multi-state fits, such as Harris:2019bih; Bali:2023sdi, multi-channel analyses through generalized eigen-value problem (GEVP) with $\pi N$ Alexandrou:2024tps; Wang:2023omf and/or $\sigma N$ states Barca:2024hrl, Feynman-Hellman inspired method Chang:2018uxx) exist, the lattice community has not yet converged on a definitive criterion for ensuring ESC is sufficiently suppressed in realistic calculations FlavourLatticeAveragingGroupFLAG:2024oxs.
@@
[…]
Table 1: Lattice spacing $a$, lattice volume $n_{L}^{3}\times n_{T}$, pion mass $m_{\pi}$ with $m_{\pi}L$, $m_{\eta_{s}}\propto\sqrt{m_{s}}$ and also the number of configurations $n_{\rm cfg}$.
aa(fm) $n_{L}^{3}\times n_{T}$ $m_{\pi}$(MeV) $m_{\pi}L$ $m_{\eta_{s}}(\rm MeV)$ $n_{\rm cfg}$
@@
Simulation Setup: The results in this work are based on the 2+1 flavor (degenerate up and down quarks plus the strange quark) ensembles CLQCD:2023sdb; CLQCD:2024yyn from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action. The information of the ensembles used in this work is summarized in Table 1.
@@
The signal-to-noise ratio of ${\cal R}_{X}$ decays exponentially with $t_{f}$, making robust control of excited-state contamination (ESC) contingent upon two factors: high-precision statistical signals at sufficiently large $t_{f}$, and the use of a proper interpolating field that suppresses contributions from excited states. The recently developed blending method Hu:2025vhd; supplemental extends the distillation method HadronSpectrum:2009krc; Morningstar:2011ka–which projects the full quark propagators onto their low-energy modes to efficiently suppress ESC–to a stochastic estimation of the full quark propagator itself. The unbiased nature of this method has been rigorously established through both mathematical proof and explicit numerical verification using 2-,3-,4-point correlation functions Hu:2025vhd, and also the nucleon iso-vector $g_{V}=0.9998(12)$ supplemental. This extension enables the precise calculation of $R_{X}(t_{f},t)$ for arbitrary combinations of $t$ and $t_{f}$ values without additional cost for quark propagator production, thereby generating the high-quality data necessary to investigate possible good choices of the interpolation operators.
@@
As detailed in the supplemental materials supplemental, our numerical tests indicate that the dominant ESC can be eliminated using a linear combination of matrix elements of ${\cal O}_{X}$ within a basis with two of interpolation fields, $H=\{{\cal N},{\cal N}_{X}\equiv{\cal N}O_{X}\}$ where the second one is “current-involved”, inspired by the discussions in Ref. Barca:2024hrl; Barca:2025det. The nucleon projection is chosen such that the matrix elements $\langle H_{i}|O_{X}|H^{\dagger}_{j}\rangle$ are non-zero for all $i,j=1,2$.
@@
Based on the studies using the heavier-than-physical light quark masses Barca:2024hrl, Ref. Barca:2025det suggested that the ESC of the correlation function $\langle{\cal N}_{X}(t_{f}){\cal O}_{X}(t){\cal N}(0)\rangle=V\langle{\cal N}(t_{f}){\cal N}(0)\rangle\langle{\cal O}_{X}(t_{f}){\cal O}_{X}(t)\rangle+{\cal O}(1)$, which contains a disconnected quark diagram, can be enhanced by the spatial volume $V$ compared to that in the standard three-point function $\langle{\cal N}(t_{f}){\cal O}_{X}(t){\cal N}(0)\rangle$. This volume-enhanced effect can be the dominant source of ESC, surpassing conventional ones. Consequently, even a small $c_{X}$ in the improved interpolating field ${\cal N}+c_{X}{\cal N}_{X}$ can significantly alter the observed ESC in the extraction of $g_{X}$. The contribution from the three-point function $\langle{\cal N}_{X}(t_{f}){\cal O}_{X}(t){\cal N}_{X}(0)\rangle$ is confirmed to be negligible, even though it entails 138 quark diagrams supplemental. This result was enabled by an automated and optimized contraction framework for arbitrary correlation functions using the blending method.
@@
In the upper panel of Fig. 1, we show the ratio ${\cal R}^{\rm mid}_{T}\equiv{\cal R}_{T}(t_{f},t_{f}/2)$ (dots) and the Feynman-Hellmann inspired combination ${\cal R}_{T}^{{\rm FH}}(t_{f})$ Chang:2018uxx (diamonds)
@@
We refer the reader to the supplemental materials supplemental for an extended discussion on the construction of the interpolation field, including its application to other typical ensembles, as well as a justification for the inefficiency of GEVP with ${\cal N}\pi$ states in the current fit range.
Results: To extract the results at the physical point in the continuum and infinite-volume limits, we employ the following joint fit ansatz:
@@
The systematic errors associated with infinite volume, continuum and chiral extrapolations are estimated by combining different fit models using the Akaike Information Criterion (AIC) BMW:2014pzb. The systematic uncertainty from ESC is estimated by taking the difference between the central values obtained from the three-state and two-state fits, and that from the renormalization constant is also integrated through the strategy used in our previous works for the quark masses CLQCD:2023sdb; CLQCD:2024yyn. The figure for the lattice spacing dependence of the results and also details on the estimate of the systematic uncertainties are provided in the Supplemental Material supplemental.
@@
In Fig. 4, we compare our results with those from the studies Bhattacharya:2015esa; Bhattacharya:2015wna; Bhattacharya:2016zcn; Gupta:2018qil; Jang:2023zts; Alexandrou:2022dtc; Harris:2019bih; Mainz24; Bali:2023sdi; QCDSFUKQCDCSSM:2023qlx; Park:2021ypf; Liu:2021irg that contribute to the current FLAG averages FlavourLatticeAveragingGroupFLAG:2024oxs. We achieve a statistical precision for $g_{T/S}$ that is improved by a factor of three or more over all previous works, with the most substantial gains on physical-point ensembles. This high precision enables a more robust and conservative evaluation of systematic errors. Our final value has a total uncertainty one-half smaller than the previous results.
Using $m_{d}-m_{u}=2.35(12)$ MeV–derived from the $N_{f}=2+1$ FLAG average FlavourLatticeAveragingGroupFLAG:2024oxs of $m_{u}/m_{d}$ and $m_{l}=(m_{u}+m_{d})/2$–and the QED correction 1.00(7)(14)-1.00(7)(14) MeV from Ref. BMW:2014pzb, we predict the neutron-proton mass difference to be Gonzalez-Alonso:2013ura:
@@
[…]
@@ Acknowledgments
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions CLQCD:2023sdb, which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. We thank Rui Zhang for valuable comments and suggestions. The calculations were performed using the PyQUDA package Jiang:2024lto with QUDA Clark:2009wm; Babich:2011np; Clark:2016rdz through HIP programming model Bi:2020wpt. The numerical calculation were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS and Advanced Computing East China Sub-center. This work is supported in part by National Key R&D Program of China No.2024YFE0109800, NSFC grants No. 12293060, 12293062, 12525504, 12435002 and 12447101, the science and education integration young faculty project of University of Chinese Academy of Sciences, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDB34030303 and YSBR-101.
@@
The blending method Hu:2025vhd splits the vector space associated with the lattice sites and color degrees of freedom, with dimension $\mathcal{L}]=N_{c}N_{L}^{3}$ into two subspaces:
1. Subspace $\mathcal{L}_{1}$ (that used in the distillation method HadronSpectrum:2009krc): The span of the $N_{\rm e}$ low-lying eigenvectors $\{v_{\lambda_{i}}\}_{i=1}^{N_{\rm e}}$ of the discrete Laplace operator;
@@
More details about the mathematical proof and numerical justification of the blending method can be found in Ref. Hu:2025vhd.
@@
Figure 5: Left: Source-sink separation $t_{f}$ dependence of $Z_{V}R_{V}(t_{f},t_{f}/2,{\cal N})_{\overrightarrow{t_{f}\rightarrow\ \infty}}g_{V}$ with different $\bar{N}_{\rm e}$, on the H48P32 ensemble, with $Z_{V}$ from the renormalization condition $Z_{V}\langle\pi^{+}|\bar{u}\gamma_{t}u|\pi^{+}\rangle=1$ of the pion matrix element of the vector current. Right: $a^{2}m_{\pi}^{2}$ dependence of the hadron-independent normalized $g_{V}$, and the result in the continuum limit is consistent with 1 within 0.12% statistical uncertainty.
@@
The left panel of Fig. 5 shows the normalized ratio $R_{V,{\cal N}}^{R}(t_{f})\equiv Z_{V}R_{V}(t_{f},t_{f}/2,{\cal N})$, which tends to $g_{V}$ as $t_{f}\rightarrow\infty$. Here, $Z_{V}\equiv 1/\langle\pi^{+}|\bar{u}\gamma_{t}u|\pi^{+}\rangle$ is the vector renormalization constant determined via the traditional sequential method CLQCD:2024yyn. As expected, the statistical uncertainty of $R_{V,\cal N}^{R}(t_{f})$ decreases with a larger $\bar{N}_{e}$, though the excited-state contamination becomes more pronounced at small $t_{f}$. A two-state fit yields $g_{V}=0.9981(9)$, which deviates from 1 by approximately $2\sigma$ with a 0.1% statistical uncertainty. This result confirms that the value of $g_{V}$ obtained with the blending method is consistent with the theoretical expectation at the 0.2% level.
@@
Since different hadron interpolation operators have different overlap with ground and excited states, the generalized eigenvalue problem (GEVP) provides a framework to suppress the low energy excited states through the combination of different interpolation operators LuscherWolff1990; Blossier:2009kd. Under GEVP, one introduces a tower of interpolation operators (GEVP basis),
@@
Then, this optimized operator can be utilized to calculate the matrix element. As shown in Ref. Blossier:2009kd, under proper choices of time slices, the ESC for the three-point functions behaves as ${\cal O}(e^{-\delta m^{\prime}\ t_{0}})$, where $\delta m^{\prime}=m_{N}-m_{0}$ is an enlarged energy gap compared to ordinary methods with a single interpolation operator in Eq. (.1). Therefore, GEVP is a large time asymptotic analysis, which systematically approaches the desired ground state with an increased number of interpolation operators.
But the question is how to choose an efficient GEVP basis for a desired matrix element. From the point of view of large time asymptotic analysis, one should start from the lowest energy states and gradually approaches the higher energy states, in the subspace constrained by the quantum number of hadron interpolation operator. Therefore, the operator list should start from the operators trying to capture the lowest energy states and proceed higher ones. For example, for nucleon matrix elements, one should start from $N$, $N\pi$, $N\pi\pi$, $N\sigma$ and so on. Recent works have explored the ESC for nucleon matrix elements under GEVP Barca:2022uhi; Grebe:2023tfx; Alexandrou:2024tin; Barca:2024hrl; Hackl:2024whw; Gao:2025loz.
@@
The concept of current-meson dominance, introduced in Refs. Barca:2024hrl; Barca:2025det, posits that certain excited states are selectively enhanced by the inserted current, becoming the dominant source of contamination in the desired matrix element. Specifically, the energy gap for this dominant state is $\Delta E=m_{X}$, where $m_{X}$ is the mass of the ground state projected by the current operator $O_{X}$. Conventional nucleon interpolation operators are inefficient at isolating this particular contamination. The solution is to include the current-involved operator ${\cal N}_{X}(\vec{0})\equiv{\cal N}(\vec{0})\sum_{\vec{z}}O_{X}(\vec{z})$, where ${\cal N}=\epsilon_{abc}u_{a}C\gamma_{5}d_{b}u_{c}$ is the standard proton interpolating field.
Consider the correlation function of ${\cal N}(t_{f})$ and ${\cal N}_{X}(0)$ with an inserted current $O_{X}(t)$, it can be decomposed into the quark line disconnected and connected diagrams,

### v3

@@ Abstract
We report a high precision calculation of the isospin vector charge $g_{S,T}$ of the nucleon using recently proposed “blending” method which provides a high-precision stochastic estimate of the all-to-all fermion propagator. Through multiplying the current operator by the traditional nucleon interpolator, we create a new operator that captures the major excited state contaminations. The linear combination of this new operator and traditional nucleon interpolator reduces these excited states and improves the robustness of the multi-state fit. Using 15 $N_{f}=2+1$ lattice ensembles which cover 5 lattice spacing, 5 combinations with the same quark masses and lattice spacing but multiple volumes, including three at the physical pion mass, we report so far most precise lattice QCD prediction $g_{T}^{\rm QCD}=1.0264[77]_{\rm tot}(53)_{\rm stat}(13)_{a}(46)_{\rm FV}(01)_{\chi}(28)_{\rm ex}(04)_{\rm re}$ and $g_{S}^{\rm QCD}=1.106[43]_{\rm tot}(31)_{\rm stat}(03)_{a}(28)_{\rm FV}(01)_{\chi}(08)_{\rm ex}(08)_{\rm re}$ at $\overline{\mathrm{MS}}$ 2 GeV, with the systematic uncertainties from continuum, infinite volume, chiral extrapolations, excited state contamination and also renormalization.
@@
The nucleon iso-vector charges $\bar{u}\Gamma_{X}u-\bar{d}\Gamma_{X}d$—specifically for the axial-vector ($\Gamma_{A}=\gamma_{5}\gamma_{i}$), scalar ($\Gamma_{S}=\mathbf{1}$), and tensor ($\Gamma_{T}=\sigma_{ij}$) currents—are among the simplest matrix elements to compute. This simplicity arises because the disconnected quark diagrams cancel out, up to isospin-breaking (ISB) effects. While the axial current can be tightly constrained from the neutron weak decay, modulo ISB and QED corrections Navas et al. (2024), the scalar and tensor charges are in higher demand. They are particularly crucial for probing TeV-scale new physics Bhattacharya et al. (2012, 2016), enabling high-precision calculations of the proton-neutron mass difference Borsanyi et al. (2015), and providing constraints on parton distribution functions Lin et al. (2018); Yao et al. (2023).
However, precise predictions of these charges remain highly challenging. This is primarily due to the competing difficulties of significant excited-state contamination (ESC) at small source-sink separations ($t_{f}$) and exponentially growing statistical noise at large $t_{f}$. Consequently, although sophisticated methods (e.g., multi-state fits, such as Harris et al. (2019); Bali et al. (2023); [LHPC 19] N. Hasan et al. (2019); [ETM 19] C. Alexandrou et al. (2020); Ottnad (2021), multi-channel analyses through generalized eigen-value problem (GEVP) with $\pi N$ Alexandrou et al. (2025); Wang et al. (2024) and/or $\sigma N$ states Barca et al. (2025), Summation Fucito et al. (1982); Capitani et al. (2012); de Divitiis et al. (2012) and Feynman-Hellman inspired method Bouchard et al. (2017); Chang et al. (2018); He et al. (2022)) exist, the lattice community has not yet converged on a definitive criterion for ensuring ESC is sufficiently suppressed in realistic calculations Aoki et al. (2024).
At the same time, a critical limitation of all existing high-precision lattice QCD results for $g_{A,S,T}$ Capitani et al. (2012); Bali et al. (2015); Bhattacharya et al. (2016); Capitani et al. (2019); Chang et al. (2018); Harris et al. (2019); Bali et al. (2023); Djukanovic et al. (2024); Jang et al. (2024) is the pion mass dependence of the finite-volume effects (FVE). One would argue that the dominance interaction at the long distance through the pion can impose $m_{\pi}^{2}$ suppression of FVE at physical quark masses Beane and Savage (2004), while it is not sufficiently constrained by the lattice data around physical pion mass which has sizable statistical uncertainty and/or specific volume. This ambiguity therefore represents an uncontrolled source of systematic error in the field.
@@
Table 1: Lattice spacing $a$, lattice volume $n_{L}^{3}\times n_{T}$, pion mass $m_{\pi}$ with $m_{\pi}L$, and also the fit values of $g_{S}$ and $g_{T}$.
aa(fm) $n_{L}^{3}\times n_{T}$ $m_{\pi}$(MeV) $m_{\pi}L$ $g_{S}$ $g_{T}$
@@
Simulation Setup: The results in this work are based on the 2+1 flavor (degenerate up and down quarks plus the strange quark) ensembles Hu et al. (2024); Du et al. (2025) from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action. The information of the ensembles used in this work is summarized in Table 1.
@@
The signal-to-noise ratio of ${\cal R}_{X}$ decays exponentially with $t_{f}$, making robust control of excited-state contamination (ESC) contingent upon two factors: high-precision statistical signals at sufficiently large $t_{f}$, and the use of a proper interpolating field that suppresses contributions from excited states. The recently developed blending method Hu et al. (2025); sup extends the distillation method Peardon et al. (2009); Morningstar et al. (2011)–which projects the full quark propagators onto their low-energy modes to efficiently suppress ESC–to a stochastic estimation of the full quark propagator itself. The unbiased nature of this method has been rigorously established through both mathematical proof and explicit numerical verification using 2-,3-,4-point correlation functions Hu et al. (2025), and also the nucleon iso-vector $g_{V}=0.9998(12)$ sup . This extension enables the precise calculation of $R_{X}(t_{f},t)$ for arbitrary combinations of $t$ and $t_{f}$ values without additional cost for quark propagator production, thereby generating the high-quality data necessary to investigate possible good choices of the interpolation operators.
@@
As detailed in the supplemental materials sup , our numerical tests indicate that the dominant ESC can be eliminated using a linear combination of matrix elements of ${\cal O}_{X}$ within a basis with two of interpolation fields, $H=\{{\cal N},{\cal N}_{X}\equiv{\cal N}O_{X}\}$ where the second one is “current-involved” which couples strongly to excited states enhanced by the current $\mathcal{O}_{X}$ itself Barca et al. (2025); Barca (2025). The nucleon projection is chosen such that the matrix elements $\langle H_{i}|O_{X}|H^{\dagger}_{j}\rangle$ are non-zero for all $i,j=1,2$.
@@
Based on the studies using the heavier-than-physical light quark masses Barca et al. (2025), Ref. Barca (2025) suggested that the ESC of the correlation function $\langle{\cal N}_{X}(t_{f}){\cal O}_{X}(t){\cal N}(0)\rangle=V\langle{\cal N}(t_{f}){\cal N}(0)\rangle\langle{\cal O}_{X}(t_{f}){\cal O}_{X}(t)\rangle+{\cal O}(1)$, which contains a disconnected quark diagram, can be enhanced by the spatial volume $V$ compared to that in the standard three-point function $\langle{\cal N}(t_{f}){\cal O}_{X}(t){\cal N}(0)\rangle$. This volume-enhanced effect can be the dominant source of ESC, surpassing conventional ones. Consequently, even a small $c_{X}$ in the improved interpolating field ${\cal N}+c_{X}{\cal N}_{X}$ can significantly alter the observed ESC in the extraction of $g_{X}$. The contribution from the three-point function $\langle{\cal N}_{X}(t_{f}){\cal O}_{X}(t){\cal N}_{X}(0)\rangle$ is confirmed to be negligible, even though it entails 138 quark diagrams sup . This result was enabled by an automated and optimized contraction framework for arbitrary correlation functions using the blending method.
@@
In the upper panel of Fig. 1, we show the ratio ${\cal R}^{\rm mid}_{T}\equiv{\cal R}_{T}(t_{f},t_{f}/2)$ (dots) and the Feynman-Hellmann inspired combination ${\cal R}_{X=T}^{{\rm FH}}(t_{f})$ Bouchard et al. (2017); Chang et al. (2018); He et al. (2022) (diamonds),
@@
We refer the reader to the supplemental materials sup for an extended discussion on the current-involved interpolator used in this work (distinct from the standard ${\cal N}\pi$ one), including its application to other typical ensembles and illustrations on the inefficiency of GEVP with ${\cal N}\pi$ interpolator in the current fit range.
Results: With significantly fewer inversions compared to prior studies [ETM 22] C. Alexandrou et al. (2023); Bali et al. (2023); Jang et al. (2024) and negligible autocorrelation sup , the values of $g_{S,T}$ obtained on all ensembles are compiled in Table 1. To extract the results at the physical point in the continuum and infinite-volume limits, we employ the following joint fit ansatz:
@@
The systematic errors associated with infinite volume, continuum and chiral extrapolations are estimated by combining different fit models using the Akaike Information Criterion (AIC) Borsanyi et al. (2015). The systematic uncertainty from ESC is estimated by taking the difference between the central values obtained from the three-state and two-state fits, and that from the renormalization constant is also integrated through the strategy used in our previous works for the quark masses Hu et al. (2024); Du et al. (2025). The figure for the lattice spacing dependence of the results and also details on the estimate of the systematic uncertainties are provided in the Supplemental Material sup .
@@
[…]
Using $m_{d}-m_{u}=2.35(12)$ MeV–derived from the $N_{f}=2+1$ FLAG average Aoki et al. (2024) of $m_{u}/m_{d}$ and $m_{l}=(m_{u}+m_{d})/2$–and the QED correction 1.00(7)(14)-1.00(7)(14) MeV from Ref. Borsanyi et al. (2015), we predict the neutron-proton mass difference [^1] In the calculation of $g_{S}$, we don’t consider $u,d$ quark mass difference. This does not affect $m_{n}-m_{p}$ at $O(m_{d}-m_{u})$ accuracy, but only contributes to $O[(m_{d}-m_{u})^{3}]$ correction that is typically ignored in the study of nucleon mass splitting effect. to be González-Alonso and Martin Camalich (2014):
@@
Conservative estimation of the FVE systematic uncertainty using the AIC with several ansatzes including both the $m_{\pi}^{2}e^{-m_{\pi}L}/\sqrt{m_{\pi}L}$ and also the naive $e^{-m_{\pi}L}$ introduces a systematic uncertainty comparable to the statistical uncertainty. This uncertainty may be universal for a wide range of hadronic matrix elements, including the non-local matrix elements required for parton distribution function (PDF) calculations within the Large Momentum Effective Theory (LaMET) framework Ji (2013); Ji et al. (2021) , and is worth follow-up studies. Consequently, previous studies of nucleon matrix elements that used the chiral suppressed ansatz may need to reassess this systematic uncertainty.
@@ Acknowledgments
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu et al. (2024), which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. We thank Rui Zhang and Srijit Paul for valuable comments and suggestions. The calculations were performed using the PyQUDA package Jiang et al. (2024) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020). The numerical calculation were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS and Advanced Computing East China Sub-center. This work is supported in part by National Key R&D Program of China No.2024YFE0109800, NSFC grants No. 12293060, 12293062, 12525504, 12435002 and 12447101, the science and education integration young faculty project of University of Chinese Academy of Sciences, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDB34030303 and YSBR-101.
@@
The blending method Hu et al. (2025) splits the vector space associated with the lattice sites and color degrees of freedom, with dimension $\mathcal{L}]=N_{c}N_{L}^{3}$ into two subspaces:
1. Subspace $\mathcal{L}_{1}$ (that used in the distillation method Peardon et al. (2009)): The span of the $N_{\rm e}$ low-lying eigenvectors $\{v_{\lambda_{i}}\}_{i=1}^{N_{\rm e}}$ of the discrete Laplace operator;
@@
More details about the mathematical proof and numerical justification of the blending method can be found in Ref. Hu et al. (2025).
@@
Figure 5: Left: Source-sink separation $t_{f}$ dependence of $Z_{V}R_{V}(t_{f},t_{f}/2,{\cal N})\ _{t_{f}\to\ \infty}\rightarrow g_{V}$ with different $\bar{N}_{\rm e}$, on the H48P32 ensemble, with $Z_{V}$ from the renormalization condition $Z_{V}\langle\pi^{+}|\bar{u}\gamma_{t}u|\pi^{+}\rangle=1$ of the pion matrix element of the vector current. Right: $a^{2}m_{\pi}^{2}$ dependence of the hadron-independent normalized $g_{V}$, and the result in the continuum limit is consistent with 1 within 0.12% statistical uncertainty.
@@
The left panel of Fig. 5 shows the normalized ratio $R_{V,{\cal N}}^{R}(t_{f})\equiv Z_{V}R_{V}(t_{f},t_{f}/2,{\cal N})$, which tends to $g_{V}$ as $t_{f}\rightarrow\infty$. Here, $Z_{V}\equiv 1/\langle\pi^{+}|\bar{u}\gamma_{t}u|\pi^{+}\rangle$ is the vector renormalization constant determined via the traditional sequential method Du et al. (2025). As expected, the statistical uncertainty of $R_{V,\cal N}^{R}(t_{f})$ decreases with a larger $\bar{N}_{e}$, though the excited-state contamination becomes more pronounced at small $t_{f}$. A two-state fit yields $g_{V}=0.9981(9)$, which deviates from 1 by approximately $2\sigma$ with a 0.1% statistical uncertainty. This result confirms that the value of $g_{V}$ obtained with the blending method is consistent with the theoretical expectation at the 0.2% level.
@@
To quantify the computational efficiency of the blending method, we compare the cost—primarily measured by the number of quark propagator inversions required to reach a target statistical precision—against traditional methods (sequential source or stochastic) used in recent high-precision studies. As a benchmark, we target similar statistical precision for $g_{S,T}$ on physical-point ensembles with $a\approx 0.08\ \mathrm{fm}$ and $L\approx 5.1\ \mathrm{fm}$.
We list here how to get the total number of inversions here:
  • PNDME, “Sequential Source Method”: $N_{\rm meas}^{\rm LP}\times(1+n_{\rm pol}\times n_{f}\times N_{\rm tsep}/N_{\rm sub})=165120\times(1+4\times 2\times 5/4)=1.8{\rm M}$
  • ETMC, “Sequential Source Method”: $\sum_{t_{\rm sep}}N_{\rm t_{sep}}^{\rm meas}\times N_{f}\times N_{\rm pol}=((750+1500+3000+4500+120000+36000+48000)\times 2\times 4=1.71{\rm M}$
  • RQCD, “Stochastic Sequential Source Method”: $n_{\rm cfg}\times(N_{\rm tsep}+N_{\rm noise}/12)=1000\times(4+100/12)=0.012{\rm M}$
  • This work, “Blending Method”: $n_{\rm cfg}\times N_{T}\times(N_{e}+N_{\rm st})/N_{c}=46\times 96\times(140+60)/3=0.39{\rm M}$
Table 2: The total inversion cost for numcleon matrix element including $g_{S}^{u-d},g_{T}^{u-d}$ among different collaboration at physical point. For a fair comparison, we list the inverse cost upto same precision of $g_{T}$ of this work in the last column.
Collaborations Ensemble L T a(fm) $m_{\pi}$(MeV) $n_{\rm cfg}$ $g_{S}$ $g_{T}$ Inversions Inversion for our precision of
$g_{S}$ $g_{T}$
ETMC(2020) cB211.072.64 64 128 0.08 139 750 1.35(17) 0.939(027) 1.71M 32M 92M
RQCD(2023) D452 64 128 0.076 156 1000 -0.6(3.0) 0.870(110) 0.01M 364M 11M
PNDME(2023) a09m130 64 96 0.09 138 1290 1.05(23) 1.010(006) 1.8M 103M 4.2M
This Work F64P13 64 128 0.078 134 46 1.00(03) 0.998(004) 0.39M 0.39M 0.39M
The comparison in Table 2 reveals our substantial efficiency gain. For five of the six comparable results from ETMC [ETM 22] C. Alexandrou et al. (2023), RQCD Bali et al. (2023), and PNDME Jang et al. (2024), we use a factor of 17 to 877 times fewer inversions to reach similar precision. The advantage is smaller but still significant (a factor of 6) for $g_{T}$ from PNDME Jang et al. (2024).
These gains stem from the method’s core design: it projects quark propagators into a low-mode subspace (like distillation) but uses stochastic techniques to estimate the full propagator within this space. This hybrid approach maximizes data reuse, allowing the calculation of N-point functions for arbitrary source-sink separations, inserted operator and hadron states, from a single set of generated “distilled” propagators.
Regarding storage cost, the primary output is these distilled propagators, which are modest in size relative to full propagators. Crucially, this same set of propagators can be reused to compute matrix elements for various hadrons and currents, amortizing the storage cost per observable and making the framework highly efficient for multi-observable campaigns.
In summary, the blending method provides a computationally and storage-efficient pathway to high-precision hadron structure calculations by achieving superior statistical precision per inversion and enabling extensive data reuse.
@@
Since different hadron interpolation operators have different overlap with ground and excited states, the generalized eigenvalue problem (GEVP) provides a framework to suppress the low energy excited states through the combination of different interpolation operators Lüscher and Wolff (1990); Blossier et al. (2009). Under GEVP, one introduces a tower of interpolation operators (GEVP basis),
@@
Then, this optimized operator can be utilized to calculate the matrix element. As shown in Ref. Blossier et al. (2009), under proper choices of time slices, the ESC for the three-point functions behaves as ${\cal O}(e^{-\delta m^{\prime}\ t_{0}})$, where $\delta m^{\prime}=m_{N}-m_{0}$ is an enlarged energy gap compared to ordinary methods with a single interpolation operator in Eq. (.1). Therefore, GEVP is a large time asymptotic analysis, which systematically approaches the desired ground state with an increased number of interpolation operators.
But the question is how to choose an efficient GEVP basis for a desired matrix element. From the point of view of large time asymptotic analysis, one should start from the lowest energy states and gradually approaches the higher energy states, in the subspace constrained by the quantum number of hadron interpolation operator. Therefore, the operator list should start from the operators trying to capture the lowest energy states and proceed higher ones. For example, for nucleon matrix elements, one should start from $N$, $N\pi$, $N\pi\pi$, $N\sigma$ and so on. Recent works have explored the ESC for nucleon matrix elements under GEVP Barca et al. (2023); Grebe and Wagman (2024); Alexandrou et al. (2024); Barca et al. (2025); Hackl and Lehner (2024); Gao et al. (2025).
@@
The concept of current-meson dominance, introduced in Refs. Barca et al. (2025); Barca (2025), posits that certain excited states are selectively enhanced by the inserted current, becoming the dominant source of contamination in the desired matrix element. Specifically, the energy gap for this dominant state is $\Delta E=m_{X}$, where $m_{X}$ is the mass of the ground state projected by the current operator $O_{X}$. Conventional nucleon interpolation operators are inefficient at isolating this particular contamination. The solution is to include the current-involved operator ${\cal N}_{X}(\vec{0})\equiv{\cal N}(\vec{0})\sum_{\vec{z}}O_{X}(\vec{z})$, where ${\cal N}=\epsilon_{abc}u_{a}C\gamma_{5}d_{b}u_{c}$ is the standard proton interpolating field.
The essential difference is that $\mathcal{N}_{X}$ yields a non-vanishing disconnected contraction, owing to the non-zero vacuum expectation value $\langle O_{X}(t)O_{X}^{\dagger}(0)\rangle$. Consider the correlation function of ${\cal N}(t_{f})$ and ${\cal N}_{X}(0)$ with an inserted current $O_{X}(t)$, it can be decomposed into the quark line disconnected and connected diagrams [^2] This equation is for the theoretical argument on the determination of current-enhanced excited states. In practical calculations, we consider all diagrams from the Wick theorem, including quark-line connected and disconnected contributions, as detailed in Sec. .2.4.,